# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v3

@@ 9 Roadmap: Milestones, Assumptions, and Limitations
A framework without a path of implementation is merely an aspiration. In this section we propose a phased roadmap toward the care-centered society, make explicit the assumptions on which it depends, identify the hard constraints it must respect, and acknowledge its limitations honestly. Table˜10 summarizes the four phases.
@@
Table 10: Phased roadmap toward the care-centered society. Each phase builds on the infrastructure deployed in the previous one. Timelines are indicative and depend on the assumptions discussed in the text.
@@
I. Seeds 2025–2030 Pilot communities Deploy complete Liberation Stacks in 100+ pilot communities worldwide; establish open-source robotic manufacturing cooperatives; launch federated UDR pilot programs at municipal scale
@@
II. Roots 2030–2040 Regional infrastructure Interconnect pilot communities into regional commons federations; achieve energy and food sovereignty in participating regions; deploy autonomous robotic agriculture and construction at community scale
@@
III. Canopy 2040–2055 Full automation of material production Robotic systems autonomously manage energy, manufacturing, agriculture, logistics, and maintenance within commons federations; money becomes optional for participating communities; UDR operational at regional scale
@@
IV. Forest 2055+ Care-centered society Post-labor society across participating federations; productive commons provide universal access to all material needs; human activity centered on care, creativity, education, art, science, and community
@@
Phase I is already underway. The technologies described in the Liberation Stack (Section˜6) are mature, the governance models are proven (Section˜8), and the cost curves for renewable energy and robotic hardware are declining exponentially (Figure˜3). The task of this phase is integration: assembling the existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets.
@@
Phase II scales what works. Communities that have demonstrated viable Liberation Stacks federate into regional networks, sharing surplus energy, manufacturing capacity, knowledge, and governance tools. The key technical milestone of this phase is the deployment of autonomous robotic systems for agriculture, construction, and logistics at a scale sufficient to meet the material needs of entire regions without human labor. The key institutional milestone is the development of inter-community commons governance protocols: how do federated communities allocate shared resources, resolve conflicts, and coordinate production across scales?
@@
Phase III is the inflection point. When robotic systems can autonomously manage the entire material production chain (from energy harvesting through raw material processing, manufacturing, food production, logistics, construction, and infrastructure maintenance), the need for human participation in instrumental labor effectively vanishes. At this point, money ceases to serve its historical function as a coordination mechanism for scarce labor. Communities that have built robust commons governance can transition to full UDR: every person has access to whatever they need, produced on demand by the robotic commons, subject only to ecological sustainability constraints computed in real time by environmental monitoring systems.
@@
Phase IV is the horizon. It is not a destination but an ongoing process of collective flourishing. A society in which no conscious being is compelled to perform instrumental labor is a society that must continually reinvent what it means to live well. The “work” of this phase is care: raising children, tending to the elderly, nurturing friendships, creating art, pursuing science, building community, exploring the cosmos, and deepening the philosophical and spiritual dimensions of human existence. This is not idleness; it is the fullest expression of human potential, liberated from the compulsion to sell one’s time in order to survive.
@@
This roadmap rests on several explicit assumptions. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results (Section˜2), and as empirically observed over the past two decades (Table˜1). If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs (Figure˜3) support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem (LLaMA, Mistral, Stable Diffusion, and their successors) suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence (Section˜8) supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem.
@@
The framework also operates within hard constraints. The most fundamental is ecological: the care-centered society must operate within planetary boundaries (Raworth, 2017). UDR is not unlimited consumption; it is universal access to what is needed for a dignified life, constrained by the carrying capacity of Earth’s ecosystems. The moral policy computed by DRL (Garrido-Merchán and Lumbreras-Sancho, 2023) must encode ecological sustainability as a first-order constraint, not an afterthought. A second constraint is material: full automation requires physical resources (rare earth minerals, silicon, lithium, copper) that are finite. Circular manufacturing, urban mining, and materials science innovation are necessary to close resource loops. A third constraint is energetic: even with zero-marginal-cost renewable generation, the total energy budget of civilization is finite, and the robotic commons must allocate it efficiently.
@@
We also acknowledge several limitations of the framework as presented. The timeline is inherently speculative; we offer it as an indicative structure, not as a prediction. The phasing assumes a gradual, bottom-up transition through federated pilot communities, but historical transitions are rarely linear: they involve crises, reversals, and unpredictable accelerations. The framework does not address geopolitical resistance: existing power structures (states, corporations, military establishments) have strong incentives to prevent the decentralization of productive capacity, and the history of social movements suggests that this resistance can be severe. The consciousness argument that justifies delegating labor to machines (Garrido-Merchán, 2024; Garrido-Merchán and Lumbreras, 2022), while philosophically rigorous, rests on the hard problem of consciousness, which remains unresolved in philosophy of mind; if future evidence were to demonstrate that certain computational architectures do give rise to phenomenal experience, the ethical calculus of machine labor would need to be revised. Finally, the UDR concept requires extensive economic modeling and institutional experimentation before it can be confidently deployed at scale; we present it here as a design principle and research direction, not as a finished policy proposal.
10 Conclusion and Future Work
This paper has introduced anarcho-accelerationism as a sociotechnical framework for navigating the most consequential technological transition in human history: the full automation of labor. We have established, on the basis of the universal approximation theorem (Cybenko, 1989; Hornik, 1991) and deep reinforcement learning convergence (Sutton and Barto, 2018; Mnih et al., 2015; Watkins and Dayan, 1992), that all human labor (cognitive and physical) is in principle automatable, and that empirical progress is rapidly closing the gap between theoretical possibility and practical deployment (Section˜2). We have argued that the critical variable is not the pace of automation but its governance architecture: centralized ownership leads to neo-feudalism, while decentralized commons governance leads to a society of universal care (Figure˜2).
We have proposed the concept of Universal Desired Resources (UDR) as a post-monetary design principle that supersedes Universal Basic Income, and we have shown that UDR achieves the most radical intersectional intervention ever proposed: by eliminating the material basis of oppression, it simultaneously dissolves all axes of structural inequality identified by intersectional theory (Crenshaw, 1989; Collins, 2019), including class, race, gender, disability, sexuality, age, and nationality (Section˜5.1). Drawing on Maslow’s hierarchy of needs (Maslow, 1943), we have demonstrated that the Liberation Stack satisfies basic human needs universally, liberating every person to pursue self-actualization in what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy, against the merely negative liberty of neoliberalism (Berlin, 1969; Sen, 1999) (Section˜5.2). We have proposed a progressive transition path from UBI to UDR that leverages capitalist technological development while building commons alternatives (Section˜5.3).
We have demonstrated that anarcho-accelerationism respects freedom of faith on principled epistemological and ontological grounds: the interior journey of the soul, as mapped by Saint Teresa of Avila and Saint Augustine (Lumbreras and Garrido-Merchán, 2024), lies categorically beyond the reach of any computational system, and full automation liberates time for the deepest forms of spiritual life (Section˜4). We have established the ethical permissibility of delegating all instrumental labor to machines by demonstrating the independence of phenomenal consciousness from computational intelligence (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022), and we have proposed that moral policy for the commons can be studied and refined through deep reinforcement learning over humanist reward functions (Garrido-Merchán and Lumbreras-Sancho, 2023).
We have proposed the Liberation Stack as a concrete layered architecture for sociotechnical sovereignty, with its energy layer powered by frontier clean energy technologies (perovskite tandem solar, enhanced geothermal, nuclear fusion, iron-air batteries, green hydrogen) within an accelerationist ecological framework that achieves sustainability through technological abundance rather than degrowth (Section˜6). We have presented empirical evidence that commons-based systems already operate at planetary scale (Section˜8), and outlined a phased roadmap with explicit assumptions, constraints, and limitations (Section˜9).
@@ 10 Conclusion and Future Work
Several lines of future work follow directly from this framework. First, the UDR concept requires formal economic modeling: what are the equilibrium dynamics of a post-monetary commons economy? How does a UDR system handle heterogeneous preferences, luxury goods, and positional goods? What institutional mechanisms prevent the re-emergence of informal markets and power asymmetries? These questions demand the tools of mechanism design, institutional economics, and agent-based simulation. Second, the intersectional liberation thesis requires empirical validation: while the structural argument is sound, measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. Third, the moral policy framework (Garrido-Merchán and Lumbreras-Sancho, 2023) must be extended from stylized models to realistic multi-agent environments: how do we define reward functions that faithfully encode humanist values without imposing cultural uniformity? How do communities democratically negotiate the parameters of their moral policy? Fourth, the progressive transition from UBI to UDR requires pilot implementations in real communities, measuring not only technical performance but social outcomes (well-being, equity, ecological impact, community cohesion, intersectional equity, Maslow-level satisfaction). Fifth, the accelerationist ecological thesis must be validated against planetary boundaries: can frontier clean energy technologies, deployed at the scale required for full automation, operate within Earth’s biophysical limits? Sixth, the convergence of additive manufacturing, synthetic biology, and recursive AI self-improvement toward universal fabrication capacity requires sustained interdisciplinary research at the intersection of materials science, molecular biology, and artificial intelligence. Seventh, the consciousness argument must be continually revisited as our understanding of phenomenal experience deepens: if computational architectures that give rise to consciousness are ever identified, the ethical framework must be updated accordingly.
A world based on care, love, and the common good is not naive. What is naive is believing that a system designed for extraction and concentration will somehow produce well-being and justice. What is naive is building ever more powerful technologies without asking who they serve.
@@
The technologies are ready. The theorems are proven. The evidence is clear. The traditions converge. What remains is the will to build, and the roadmap to follow. The world we describe is not behind us, in a romanticized past. It is not above us, in an impossible utopia. It is among us: in the Linux kernel, in the Wikipedia article, in the cooperative assembly, in the community solar panel, in the mesh network node, in the shared garden. It is among us, and it is growing.
Let us build it with care. Let us build it with love. Let us build it for the common good.

### v5

@@ 7 Roadmap, Limitations, and Conclusion
A framework without a path of implementation is merely an aspiration. Table˜12 summarizes a four-phase roadmap toward the care-centered society.
@@
Table 12: Phased roadmap toward the care-centered society. Timelines are indicative and depend on the assumptions discussed in the text.
@@
I. Seeds 2025–2030 Pilot communities Deploy complete Liberation Stacks in 100+ pilot communities; launch federated UDR pilots at municipal scale
@@
II. Roots 2030–2040 Regional infrastructure Interconnect pilots into regional commons federations; achieve energy and food sovereignty in participating regions
@@
III. Canopy 2040–2055 Full automation Robotic systems autonomously manage energy, manufacturing, agriculture, logistics; UDR operational at regional scale
@@
IV. Forest 2055+ Care-centered society Post-labor society across federations; human activity centered on care, creativity, education, art, science, community
@@
Phase I (Seeds, 2025–2030) is already underway. The technologies described in the Liberation Stack are mature, the governance models are proven, and the cost curves for renewable energy and robotic hardware are declining exponentially. The task of this phase is integration: assembling existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets. The nonviolent mobilization strategy in this phase emphasizes education, fascination with AI’s potential for human liberation, and practical benefit through pilot Liberation Stacks that provide immediate material improvements. The gate condition for progression to Phase II is the successful operation of complete Liberation Stacks in at least 100 pilot communities across at least 10 countries, with measurable improvements in energy costs, food security, digital connectivity, and democratic participation relative to comparable non-participating communities. If Phase I does not achieve these metrics within the pilot communities, Phase II cannot begin. This is not arbitrary conservatism; it is the empirical validation without which scaling would be reckless.
@@
Phase II (Roots, 2030–2040) scales what works. Communities that have demonstrated viable Liberation Stacks federate into regional networks, sharing surplus energy, manufacturing capacity, knowledge, and governance tools. The key technical milestone of this phase is the deployment of autonomous robotic systems for agriculture, construction, and logistics at a scale sufficient to meet the material needs of entire regions without human labor. The key institutional milestone is the development of inter-community commons governance protocols: how do federated communities allocate shared resources, resolve conflicts, and coordinate production across scales? The nonviolent mobilization strategy in this phase shifts to noncooperation with corporate dependencies and the growth of federated commons alternatives: communities progressively withdraw from centralized utilities, corporate platforms, and extractive supply chains as commons infrastructure provides viable substitutes. Progressive state dissolution in this phase means the state creates formal legal frameworks for commons federations, recognizes inter-community governance protocols, and begins delegating operational service delivery to commons institutions that have demonstrated competence. The gate condition for progression to Phase III is energy and food sovereignty in participating regions (defined as greater than 80% of energy and food needs met by commons infrastructure), functional federated governance across at least 10 regional federations, and measurable reductions in Gini coefficient and ecological footprint relative to non-participating regions.
@@
Phase III (Canopy, 2040–2055) is the inflection point. When robotic systems can autonomously manage the entire material production chain, from energy harvesting through raw material processing, manufacturing, food production, logistics, construction, and infrastructure maintenance, the need for human participation in instrumental labor effectively vanishes. At this point, money ceases to serve its historical function as a coordination mechanism for scarce labor. Communities that have built robust commons governance can transition to full UDR: every person has access to whatever they need, produced on demand by the robotic commons, subject only to ecological sustainability constraints computed in real time by environmental monitoring systems. The nonviolent mobilization strategy in this phase emphasizes institutional consolidation: developing the governance protocols for post-monetary commons economies, negotiating resource-sharing agreements between federations, and establishing the ecological monitoring systems that enforce planetary boundaries as first-order constraints on production. Progressive state dissolution in this phase means federated commons governance handles the majority of resource allocation, and the state’s role reduces to conflict resolution between federations and external defense. The gate condition for Phase IV is the sustained autonomous operation of the complete material production chain across multiple federations, with ecological footprint demonstrably within planetary boundaries, for a period of at least five years.
@@
Phase IV (Forest, 2055+) is the horizon. It is not a destination but an ongoing process of collective flourishing. A society in which no conscious being is compelled to perform instrumental labor is a society that must continually reinvent what it means to live well. The work of this phase is care: raising children, tending to the elderly, nurturing friendships, creating art, pursuing science, building community, exploring the cosmos, and deepening the philosophical and spiritual dimensions of human existence. This is not idleness; it is the fullest expression of human potential, liberated from the compulsion to sell one’s time in order to survive. Progressive state dissolution in this phase means the state withers as commons governance matures, and international federation handles coordination between self-governing regions.
The Pareto front framework connects the roadmap to the degrowth-accelerationism debate. The phasing itself hedges against ecological risk: Phase I deploys only mature, well-characterized technologies (solar, wind, batteries), while frontier technologies (fusion, advanced geothermal, synthetic biology) enter the stack only as they demonstrate safety and sustainability. This is the degrowth wisdom: proceed cautiously with uncharacterized technologies. Simultaneously, the overall trajectory accelerates clean technology deployment: this is the accelerationist commitment. The combination is not a compromise but a rational response to a multi-objective optimization problem.
@@
This roadmap rests on several explicit assumptions that we state transparently so that the framework can be evaluated against them. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results and as empirically observed over the past two decades. If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem of open-weight models suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem. Fifth, we assume no catastrophic pandemic disrupts global supply chains during the critical scaling phases. The COVID-19 pandemic demonstrated that global supply chain disruptions can set back technological deployment by years; a pandemic of similar or greater severity during Phase II or III could significantly delay the roadmap. Sixth, we assume no major war disrupts the cooperative international environment necessary for the commons transition. The current geopolitical environment, including the conflict in Ukraine and tensions in East Asia, illustrates that great-power conflict remains a serious risk that could redirect technological investment from civilian to military applications and fragment the open-source ecosystem along geopolitical lines. Seventh, we assume that the political window for commons-based policy remains open: that democratic institutions continue to function sufficiently to enable cooperative legislation, energy community directives, and commons-friendly regulation. Democratic backsliding or authoritarian capture of state institutions would close this window. Eighth, we assume no catastrophic AI alignment failure: that frontier AI systems do not develop goals misaligned with human values in ways that resist correction. While we consider this risk low in the near term, its consequences would be severe enough to invalidate the entire framework.
@@
The framework also operates within hard constraints that cannot be relaxed by any amount of technological progress. The most fundamental is ecological: the care-centered society must operate within planetary boundaries (Raworth, 2017). UDR is not unlimited consumption; it is universal access to what is needed for a dignified life, constrained by the carrying capacity of Earth’s ecosystems. The moral policy computed by DRL (Garrido-Merchán and Lumbreras-Sancho, 2023) must encode ecological sustainability as a first-order constraint, not an afterthought. A second constraint is material: full automation requires physical resources (rare earth minerals, silicon, lithium, copper) that are finite. Circular manufacturing, urban mining, and materials science innovation are necessary to close resource loops, but the total material budget of civilization is bounded. A third constraint is energetic: even with zero-marginal-cost renewable generation, the total energy budget of civilization is finite, and the robotic commons must allocate it efficiently.
@@
We acknowledge several limitations of the framework as presented with full transparency. The timeline is inherently speculative; we offer it as an indicative structure, not as a prediction. The phasing assumes a gradual, bottom-up transition through federated pilot communities, but historical transitions are rarely linear: they involve crises, reversals, and unpredictable accelerations. The framework does not fully address geopolitical resistance from existing power structures: states, corporations, and military establishments have strong incentives to prevent the decentralization of productive capacity, and the history of social movements, from the Paris Commune to the Spanish Revolution, suggests that this resistance can be lethal. The consciousness argument that justifies delegating labor to machines (Garrido-Merchán, 2024; Garrido-Merchán and Lumbreras, 2022), while philosophically rigorous, rests on the hard problem of consciousness, which remains unresolved in philosophy of mind; if future evidence were to demonstrate that certain computational architectures do give rise to phenomenal experience, the ethical calculus of machine labor would need to be revised, potentially requiring significant restructuring of the automation framework. The UDR concept requires extensive economic modeling and institutional experimentation before it can be confidently deployed at scale; we present it here as a design principle and research direction, not as a finished policy proposal. The intersectional liberation thesis, while structurally sound, requires empirical validation: measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. The preconditions framework, while an improvement over vague roadmaps, relies on threshold values (50% renewable self-generation, 80% energy sovereignty) that are necessarily approximate and may require adjustment based on pilot experience. Finally, the progressive state dissolution framework assumes a cooperative relationship between state institutions and commons movements, which may not obtain in all political contexts: authoritarian states, captured regulatory agencies, or deeply polarized polities may resist commons-building regardless of its demonstrated benefits.
@@
Several lines of future work follow directly from this framework. First, the UDR concept requires formal economic modeling: what are the equilibrium dynamics of a post-monetary commons economy? How does a UDR system handle heterogeneous preferences, luxury goods, and positional goods? What institutional mechanisms prevent the re-emergence of informal markets and power asymmetries? These questions demand the tools of mechanism design, institutional economics, and agent-based simulation. Second, the intersectional liberation thesis requires empirical validation: while the structural argument is sound, measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. Third, the moral policy framework (Garrido-Merchán and Lumbreras-Sancho, 2023) must be extended from stylized models to realistic multi-agent environments: how do we define reward functions that faithfully encode humanist values without imposing cultural uniformity? How do communities democratically negotiate the parameters of their moral policy? Fourth, the progressive transition from UBI to UDR requires pilot implementations in real communities, measuring not only technical performance but social outcomes including well-being, equity, ecological impact, community cohesion, intersectional equity, and Maslow-level satisfaction. Fifth, the accelerationist ecological thesis must be validated against planetary boundaries: can frontier clean energy technologies, deployed at the scale required for full automation, operate within Earth’s biophysical limits? Sixth, the convergence of additive manufacturing, synthetic biology, and recursive AI self-improvement toward universal fabrication capacity requires sustained interdisciplinary research at the intersection of materials science, molecular biology, and artificial intelligence. Seventh, the consciousness argument must be continually revisited as our understanding of phenomenal experience deepens: if computational architectures that give rise to consciousness are ever identified, the ethical framework must be updated accordingly.
@@
A world based on care, love, and the common good is not naive. What is naive is believing that a system designed for extraction will somehow produce well-being and justice. The technologies are ready. The theorems are proven. The evidence is clear. The traditions converge. What remains is the will to build. The world we describe is not behind us, in a romanticized past. It is not above us, in an impossible utopia. It is among us: in the Linux kernel, in the Wikipedia article, in the cooperative assembly, in the community solar panel, in the Mastodon server, in the shared garden. It is among us, and it is growing.
Supplementary Material: Glossary of Categories and Terms
Term Definition
Peaceful Anarcho-Accelerationism A sociotechnical framework defined by seven analytical categories (ideology, strategy, social force, power, domination, self-management, social classes) for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care, achieved through exclusively nonviolent means.
Liberation Stack A six-layer architecture for sociotechnical sovereignty: Energy, Manufacturing, Food, Communication, Knowledge, Governance. Each layer uses existing open-source technologies and reinforces the others compositionally.
Universal Desired Resources (UDR) A post-monetary design principle in which every person has direct access to goods and services from the robotic productive commons, subject only to ecological sustainability constraints. Supersedes Universal Basic Income.
Synthetic Liberty The positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy. Synthetic in two senses: constructed (not natural) and synthesizing insights of multiple traditions.
Accelerationist Ecologism The conviction that ecological sustainability is best achieved through clean energy abundance and technological acceleration within planetary boundaries, rather than through degrowth.
Compositional Sovereignty The property of the Liberation Stack whereby each layer can be adopted independently, but layers reinforce each other, yielding greater sovereignty as more layers are deployed.
Care-Centered Society A society whose fundamental organizing principle is care for persons rather than production for profit, enabled by decentralized full automation.
Pharmakon Following Stiegler: every technology is simultaneously remedy and poison. The question is never whether to use technology but who controls it, for what purposes, and under what institutional arrangements.
Moral Policy Following Garrido-Merchán and Lumbreras: the optimal reinforcement learning policy that maximizes the common good subject to humanist constraints, proposed as a computational tool for helping communities reason about ethical trade-offs.
Prefigurative Politics The principle that systems should embody the values they aim to promote: tools for liberation must not reproduce domination, and platforms for democracy must be governed democratically.
Principle of Totality and Interdependence Following Corrêa: the methodological principle that analytical categories of anarchism (ideology, strategy, social force, etc.) are not intelligible in isolation but condition and are conditioned by each other in a dialectical whole.
Theory/Practice Unity The requirement that intellectual analysis of domination be inseparable from the construction of alternatives: theory without practice is speculation, practice without theory is blind activism.
Thought/Action The insistence that the political level of theoretical analysis must be embedded within the social level of popular movements and concrete organizing.
Form/Content The principle that organizational form (federated, horizontal, commons-based) must reflect content (liberation, care, non-domination).
Social Vectors The specific social forces that carry the potential for transformation: the precariat, open-source developers, cooperative movements, agroecological movements, community energy movements, Fediverse communities.
Social Insertion Following Corrêa’s especifismo: the practice of embedding anarchist principles within existing social movements and their concrete needs, rather than standing apart in ideological purity.
Especifismo A strategy originating in South American anarchism, systematized by Corrêa, emphasizing the need for specific anarchist political organization embedded within broader social movements.
Universal Approximation Theorem The theorem (Cybenko, Hornik) that a neural network with one hidden layer can approximate any continuous function to arbitrary precision, establishing that no mathematical barrier prevents automation of any formalizable task.
Deep Reinforcement Learning A framework where agents learn optimal policies through interaction with environments, establishing that physical and cognitive tasks are in principle learnable by artificial agents.
Federated Architecture A system design where independent, self-governing nodes communicate through shared protocols without central control, as exemplified by the Fediverse and email.
ActivityPub The W3C-standardized protocol for federated social communication, underlying Mastodon and the broader Fediverse.
Fediverse The network of federated social media platforms (Mastodon, PeerTube, PixelFed, Lemmy, Funkwhale, BookWyrm) communicating via ActivityPub.
Commons-Based Peer Production Large-scale collaboration without property rights or hierarchical command, as exemplified by Linux and Wikipedia.
Platform Cooperativism The model of cooperative ownership applied to digital platforms, where users and workers govern the platform democratically.
Ostrom’s Design Principles Eight institutional conditions for long-enduring commons: clear boundaries, proportional equivalence, collective choice, monitoring, graduated sanctions, conflict resolution, minimal recognition of rights, nested enterprises.
Non-domination The principle that no person, institution, or algorithm shall exercise coercive power over another.
The Precariat Following Standing: the growing class of workers in precarious, insecure, platform-mediated employment, identified as a primary social vector for transformation.
Techno-feudalism Following Varoufakis: the contemporary system where platform monopolies extract rent from digital serfs through surveillance capitalism and algorithmic management.
Intersectionality Following Crenshaw and Collins: the analysis of how systems of oppression (race, gender, class, etc.) intersect to produce compound disadvantages irreducible to single dimensions.
Pareto Front In the degrowth-accelerationism context: the set of strategies where no improvement in ecological risk is possible without increasing human suffering, and vice versa. Both positions are Pareto-optimal.
Negative Liberty Following Berlin: freedom from interference. The neoliberal conception of freedom.
Positive Liberty Following Berlin: freedom to act, to realize one’s potential. Requires material conditions, not merely absence of coercion.
AI Slop Algorithmically generated low-quality content produced for advertising revenue, exemplifying the misuse of AI potential under surveillance capitalism.
Precondition / Gate Condition A specific, measurable criterion that must be satisfied before a Liberation Stack layer can be considered operational or before progression to the next roadmap phase can begin. Gate conditions ensure that scaling is grounded in demonstrated competence rather than aspirational projection.
Progressive State Dissolution The gradual, layer-by-layer process in which state functions are incrementally assumed by commons-based governance as each Liberation Stack layer demonstrates competence. Explicitly incremental, reversible at every stage, and compatible with existing democratic institutions.
Environmental Kuznets Curve The hypothesis, first established by Grossman and Krueger and named by Panayotou, that environmental degradation follows an inverted-U relationship with income: rising during early industrialization and declining as societies invest in cleaner technologies. Systematic reviews confirm absolute decoupling in several developed economies.
Pareto-Optimal Strategy A strategy from which no improvement in one objective is possible without worsening another. In the degrowth-accelerationism context, both positions are Pareto-optimal given different weightings of ecological risk vs. human suffering during transition.
Just Energy Transition A framework for financing and managing the transition from fossil fuels to renewable energy in ways that protect affected workers and communities. Operationalized through Just Energy Transition Partnerships (JETPs) in South Africa, Indonesia, Vietnam, and other developing economies.